# An improved description of nuclear charge radii: Global trends beyond $N=28$ shell closure

### v1

@@ Abstract
Charge radii measured with high accuracy provide a stringent benchmark for characterizing nuclear structure phenomena. In this work, the systematic evolution of charge radii for nuclei with $Z=19$-$29$ is investigated through relativistic mean field (RMF) theory with effective forces NL3, PK1 and $*{}^{*}$. The neutron-proton ($np$) correlation around Fermi surface originated from the unpaired neutron and proton has been taken into account tentatively in order to reduce the overestimated odd-even staggering (OES) of charge radii. This improved method can give an available description of charge radii across $N=28$ shell closure. A remarkable observation is that the charge radii beyond $N=28$ shell closure follow the similarly steep increasing trend, namely irrespective of the number of protons in the nucleus. Especially, the latest results of charge radii for nickel and copper isotopes can be reproduced remarkably well. Along $N=28$ isotonic chain, the sudden increase of charge radii is weakened across $Z=20$, but presented evidently across $Z=28$ closed shell. The abrupt changes of charge radii across $Z=22$ are also shown along $N=32$ and $34$ isotones, but latter with a less slope. This seems to provide a sensitive indicator to identify the new magicity of a nucleus with universal trend of charge radii.
@@ I INTRODUCTION
Nuclear charge radius encodes the information about the charge density distribution in a self-bound many-nucleon system. Reliable description of nuclear charge radius can provide access to recognize various pronounced structure phenomena, such as halo structures Geithner et al. (2008); Nörtershäuser et al. (2009), shape staggering Marsh et al. (2018); Barzakh et al. (2017); Sels et al. (2019); Barzakh et al. (2021), the onset of shape deformation Lalazissis et al. (1996); Casten (1985); Togashi et al. (2016) and the emergence of nuclear magicity Kreim et al. (2014); Gorges et al. (2019); Novario et al. (2020); Bagchi et al. (2019); Kaur et al. (2022), etc. Meanwhile, charge radii difference of mirror-pair nuclei can provide a probe to constrain the isovector components of equation of state (EoS) in asymmetric nuclear matter Wang and Li (2013); Brown (2017); Yang and Piekarewicz (2018); Brown et al. (2020); Xu et al. (2022); Pineda et al. (2021); Novario et al. (2023); Huang et al. (2023); An et al. (2023); Bano et al. (2023); König et al. (2023a). With the improved measurements, much more data of charge radii of exotic nuclei far away from the $\beta$-stability line have been complied Angeli and Marinova (2013); Li et al. (2021); Day Goodacre et al. (2021); Yang et al. (2023). The distinguished aspects of nuclear charge radii, involving the shell closures and odd-even staggering (OES) effects, are generally observed throughout the entire nuclide chart Campbell et al. (2016); Garcia Ruiz and Vernon (2020).
@@
The discontinuity changes in charge radii are naturally observed across the traditional neutron magic numbers $N=28$, $50$, $82$ and $126$ Procter et al. (2012); Gorges et al. (2019); Sharma et al. (1993); Goddard et al. (2013); Hammen et al. (2018); de Groote et al. (2020). Many methods pertaining to the effect of shell closure on nuclear charge radii are discussed Wang and Li (2013); Nakada (2019); Perera et al. (2021). As demonstrated in Ref. Sharma et al. (1995) that this shell closure effect results from the rather small isospin dependence of the spin-orbit term. Especially striking among these fully filled shells has been found in the intermediate mass region nuclei with featuring the $N=28$ neutron-closure shell. The inverted parabolic-like shape between $N=20$ and $N=28$ and OES in charge radii are observed remarkably along Ca isotopes. The same scenario can also be found in K isotopes, but with weakened amplitudes Angeli and Marinova (2013). Across the $N=28$ shell closure, the experimentally observed strong increases in charge radii are naturally manifested for K Koszorús et al. (2019, 2021) and Ca Garcia Ruiz et al. (2016) isotopes, and the odd-even oscillation behaviors have been weakened as well. Recently, charge radii of ${}^{55,56,58-68,70}\mathrm{Ni}$ isotopes have been measured by collinear laser spectroscopy method Malbrunot-Ettenauer et al. (2022); Sommer et al. (2022). It suggests that the trend of charge radii across the $N=28$ shell closure is similar to Ca isotopes. Between neutron numbers $N=28$ and $N=40$, the nuclear charge radii exhibit a universal pattern that is independent of atomic number Kortelainen et al. (2022). These local variations pose great challenges to our knowledge of understanding nuclear force.
@@
The undertaken efforts have been devoted to describing the global trend of charge radii in nuclear chart. The sophisticated algebraic expression involving the Casten factor $P$ can also reproduce the shell effects Casten et al. (1987); Angeli (1991). Subsequently, a five-parameter formula has been proposed by introducing the OES effects Sheng et al. (2015). In which neutron-proton ($np$) interaction originating from the valence proton and neutron contributes to the local variation of nuclear size. This indicates that the $np$-correlation plays an essential role in accessing the fine structure of nuclear size. Miller, $et~{}al.$ point out that short-range neutron-proton tensor interactions cause the protons to move closer to the outside neutrons of a nucleus, thereby increasing the charge radius Miller et al. (2019a). This emphasizes that $np$-correlation results in the changes of the calculated root-mean-square (rms) charge radius. This short-range correlations (SRCs) imply a decrease in the occupation of the states below the Fermi level and a partial occupation in the states above it Souza et al. (2020). Recent study further verifies that the nuclear SRCs induce the high-momentum fluctuations in the nuclear medium, and the role of $np$-correlation has a non-negligible influence on nuclear size Cosyn and Ryckebusch (2021).
@@
As demonstrated in Ref. Zawischa (1985), there is an attractive interaction between proton and neutron pairs, the origin of OES behaviors in nuclear charge radii is due to the four-particle correlations (or being $\alpha$-cluster). The relativistic density functionals cannot describe the fine structure of charge radii along calcium isotopic chain well Geng et al. (2003); Blaum et al. (2008); Xia et al. (2018); Zhang et al. (2022). To reproduce the OES and shell effect of charge radii along calcium isotopes, a modified relativistic mean field plus BCS (RMF(BCS)*) ansatz has been proposed An et al. (2020). In which the semi-microscopic correction originating from the neutron and proton pairs condensation is introduced by solving the state-dependent BCS equations. The calculated results are in good agreement with the data and compatible with the sophisticated Fayans energy density functional (EDF) approach where the surface pairing dominates a role in reproducing the local variation of charge radii Reinhard and Nazarewicz (2017); Miller et al. (2019b). However, for odd-$Z$ isotopic chain, the OES effects in charge radii are overestimated by RMF(BCS)* approach An et al. (2022a). This may result from the fact that the correlation between the unpaired neutron and proton due to the violation of time-reversal symmetry is ignored. The same scenarios can also be encountered in Fayans EDF model where the OES effects in charge radii are overestimated for K Koszorús et al. (2021) and Cu de Groote et al. (2020) isotopes as well. Obviously, the influence coming from the unpaired nucleons is still excluded out the pairing spaces due to the blocking approximation.
@@
So far, the existing theoretical models can hardly give a unified description of charge radii around $N=28$ mass region. Hence the required model should be able to describe simultaneously the shell effect and OES in nuclear charge radii. In this work, the charge radii are calculated along $Z=19$-$29$ isotopic chains that include the nuclei featuring the $N=28$ magic shell. In order to reduce the overestimated OES behaviors of charge radii along odd-$Z$ isotopic chains, the $np$-correlation originating from the simultaneously unpaired neutron and proton is tentatively taken into account in the improved charge radii formula. One should also mention the great interest for the properties of the neutron number $N=32$ and $34$ isotopes that have been proposed as new-magicity nuclei in certain isotopic chains. The sensitive indicators of new-magicity associating with the sudden increase of charge radii are also paid more attention along $N=32$ and $34$ isotonic chains.
@@ II Theoretical framework
The relativistic mean field (RMF) theory with different parametric versions has made considerable success in describing various phenomena in nuclear physics Meng and Ring (1996); Vretenar et al. (2005); Liang et al. (2015); Meng (2016); Cao and Ma (2004); Paar et al. (2003, 2004); Zhou et al. (2003); Meng et al. (2006); Zhou et al. (2010); Li et al. (2012); Xu et al. (2015); Niks̆ić et al. (2014); Sun et al. (2018); Zhang et al. (2020, 2021); Rong (2023). In this work, the standard non-linear self-coupling Lagrangian density is employed Ring et al. (1997). The Dirac equations with effective fields $S(\mathbf{r})$ and $V(\mathbf{r})$ are derived through variational principle. To obtain the ground state properties of finite nuclei, the quadrupole deformation parameter $\beta_{20}$ becomes constraint in the self-consistent iterative process. Therefore, the Hamiltonian formalism can be recalled as follows
@@
where $M$ is the mass of nucleon, $\lambda$ is the spring constant and $\mathbf{Q}$ is the intrinsic quadrupole moment. Those values of $\beta_{20}$ are changed from $-0.50$ to $0.50$ with the interval range of $0.01$.
[…]
@@
The first term represents the charge distribution of point-like protons and the second term is due to the finite size of protons Gambhir et al. (1990). Here, the quantity of the proton radius takes the values about $0.84$ fm Tiesinga et al. (2021); Lin et al. (2022). As shown in Eq. (2), the quantity of $\mathcal{D}$ is associated with the Cooper pairs condensation Dussel et al. (2007). Furthermore, this quantity can be used to measure the Fermi surface ‘diffuseness’ encoded by various eigenfunctions Mang et al. (1965). Its value is calculated self-consistently by solving the state-dependent BCS equations Geng et al. (2003); An et al. (2018). The difference of $\mathcal{D}$ coming from neutrons and protons has been attributed to the $np$-correlation around Fermi surface An et al. (2020, 2022b).
The last term in this expression just represents the $np$-correlation deriving from the simultaneously unpaired neutron and proton. This means the quantity of $\delta$ equals to zero for even-even, odd-even and even-odd nuclei. The values of $a_{0}=0.561$ and $\delta=0.355$ are calibrated by fitting the odd-even oscillation and the inverted parabolic-like shape of charge radii along potassium and calcium isotopes under effective force NL3. And then the extrapolated calculations are performed for $Z=21$-$29$ isotopic chains. The PK1 set gives a more reasonable baryonic saturation density and a reasonable description of spin-orbit splittings and single-particle energies Long et al. (2004). In addition, $*{}^{*}$ parameter set improves the description of the ground state properties of many nuclei and simultaneously provides an excellent description of the collective excited states properties in finite nuclei Lalazissis et al. (2009). Thus these two effective forces are also employed to access the global evolution of charge radii along $Z=19$-$29$ isotopic chains. It should be particularly interesting for quantitatively understanding the reduced OES of charge radii along odd-$Z$ isotopic chains and the trend of changes of nuclear charge radii across the $N=28$ shell closure.
@@ III Results and Discussions
[…]
@@ III.1 Shell quenching of nuclear charge radii at $N=28$
The inverted parabolic-like behavior describing the charge radii of the calcium isotopes between $N=20$ and $28$ as well as the amplitude of OES is much more pronounced for the neighbouring odd-$Z$ element potassium Touchard et al. (1982). The inverted parabolic-like curves are symmetric to the $1f_{7/2}$ mid-shell neutron number $N=24$ for potassium and calcium isotopic chains Angeli and Marinova (2013). In Fig. 1, the rms charge radii of potassium and calcium isotopes as a function of neutron number are plotted by the relativistic mean field (RMF) model with effective forces NL3, PK1 and $*{}^{*}$. One can find that the inverted parabolic-like shapes of potassium and calcium isotopes are compared against well by NL3 and $*{}^{*}$ parametrization sets, while the quantitative level of charge radii is systematically underestimated by PK1 force. Beyond $N=28$ shell closure, the sharp increases are identified by these three forces. In Ref. Borzov and Tolokonnikov (2022), the universal increase of the charge radii after crossing the neutron shell at $N=28$ is attributed to the quasiparticle-phonon coupling.
@@
Figure 1: Evolution in rms charge radii versus neutron number for potassium (a) and calcium (b) isotopes. The effective forces NL3 (open circle), PK1 (open square) and $*{}^{*}$ (open triangle) are employed in relativistic mean field framework. The corresponding odd-even staggerings are also depicted in panels (c) and (d). Experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021), which are represented by solid diamond with systematic error bands (light blue).
@@
The characteristic aspects in charge radii are the odd-even oscillation behaviors along calcium isotopic. Ref. Talmi (1984) demonstrates that the main contribution to the odd-even effect is ascribed to quadrupole and higher even-order core polarization. In which the strong and attractive interaction between the valence neutrons and core protons has been considered. These peculiarities have been ascribed to the changes of dynamic deformation for different orders Barranco and Broglia (1985). Qualitatively the OES of radii as a function of $N$ is associated with the reduction of core polarization due to unpaired nucleons Reehal and Sorensen (1971). This mechanism also provides a straightforward explanation for the odd-even effect as a function of $Z$. Besides, high-order radial moment offers key information on the nuclear charge radius Reinhard et al. (2020).
@@
As shown in Ref. Koszorús et al. (2021), the amplitudes of OES in charge radii for potassium isotopes are overestimated by Fayans EDF. Fayans EDF model points out that surface pairing interactions cause the origin of OES in nuclear charge radii Reinhard and Nazarewicz (2017). However, the time-reversal states are excluded in tackling pairing correlations, especially for simultaneously unpaired neutron and proton in potassium isotopes. The same scenario is encountered in Ref. An et al. (2022a), where the states that violate time-reversal symmetry are tackled by the blocking approximation. This leads to the most overestimated amplitudes of OES of charge radii in potassium isotopes as well. In Ref. An et al. (2020), $np$-correlation originating from the difference of Cooper pair condensation is introduced by solving the state-dependent BCS equations. The modified root mean square charge radii formula can describe well the OES and shell effects in charge radii along even-$Z$ isotopic chains. But for odd-$Z$ cases, the OES is overestimated due to the absence of $np$-correlation coming from the simultaneously unpaired neutron and proton An et al. (2022a). As shown in Fig 1, the modified formula can reproduce the OES of charge radii between $N=20$ and $28$ along potassium and calcium isotopes.
@@
In Fig. 1 (a), charge radii of ${}^{36-38}$ isotopes are slightly overestimated by the NL3 and $*{}^{*}$ forces. By contrast, the calculated results for ${}^{50}$ are slightly underestimated by these two forces. The experimental mass measurement suggests that ${}^{51}$ has a relative large shell gap with respect to the neighbouring counterparts Rosenbusch et al. (2015). Of high interest is also the experimental determination of the charge radii across the neutron number $N=32$ with emerging the new magicity. A signature of the magic character can be reflected from the sudden increase in charge radii. Charge radii of exotic isotope ${}^{52}$ has been performed by the collinear resonance ionization spectroscopy method Koszorús et al. (2021). This latest measurement suggests that no rapid increase in charge radii can be found across ${}^{51}$. Beyond ${}^{50}$, these three effective forces give almost similar trends and show a smooth increase from ${}^{51}$ to ${}^{52}$.
@@
In Fig. 1 (b), charge radii of ${}^{37-39}$a isotopes obtained by PK1 force are compared well against NL3 and $*{}^{*}$ forces. $*{}^{*}$ parameter set gives almost comparable results with respect to NL3 force. With the increasing neutron number occupation at the $p_{3/2}$ orbital, the unexpected linear increase of the charge radius occurs Enciu et al. (2022). This can also be used to understand the emergency of the abrupt increase of charge radii across $N=28$ shell closure as shown in Fig. 1. The neutron number $N=32$ is identified as a new magic number in calcium region An et al. (2018); Wienholtz et al. (2013); Gallant et al. (2012). Therefore, the implication of this experimental value for the charge radius of ${}^{53}$a will be urgently expected. As encountered in potassium isotopes, the charge radii beyond $N=34$ show same steep increasing trend with these effective forces.

### v2

# Improved description of nuclear charge radii: Global trends beyond $N=28$ shell closure

@@ Abstract
Charge radii measured with high accuracy provide a stringent benchmark for characterizing nuclear structure phenomena. In this work, the systematic evolution of charge radii for nuclei with $Z=19$-$29$ is investigated through relativistic mean field theory with effective forces NL3, PK1, and NL3. The neutron-proton ($np$) correlation around Fermi surface originated from the unpaired neutron and proton has been taken into account tentatively in order to reduce the overestimated odd-even staggering of charge radii. This improved method can give an available description of charge radii across $N=28$ shell closure. A remarkable observation is that the charge radii beyond $N=28$ shell closure follow the similarly steep increasing trend, namely irrespective of the number of protons in the nucleus. Especially, the latest results of charge radii for nickel and copper isotopes can be reproduced remarkably well. Along $N=28$ isotonic chain, the sudden increase of charge radii is weakened across $Z=20$, but presented evidently across $Z=28$ closed shell. The abrupt changes of charge radii across $Z=22$ are also shown along $N=32$ and $34$ isotones, but the latter with a less slope. This seems to provide a sensitive indicator to identify the new magicity of a nucleus with universal trend of charge radii.
@@ I INTRODUCTION
Nuclear charge radius encodes the information about the charge density distribution in a self-bound many-nucleon system. Reliable description of nuclear charge radius can provide access to recognize various pronounced structure phenomena, such as halo structures Geithner et al. (2008); Nörtershäuser et al. (2009), shape staggering Marsh et al. (2018); Barzakh et al. (2017); Sels et al. (2019); Barzakh et al. (2021), the onset of shape deformation Lalazissis et al. (1996); Casten (1985); Togashi et al. (2016), and the emergence of nuclear magicity Kreim et al. (2014); Gorges et al. (2019); Novario et al. (2020); Bagchi et al. (2019); Kaur et al. (2022), etc. Meanwhile, charge radii difference of mirror-pair nuclei can provide a probe to constrain the isovector components of equation of state (EoS) in asymmetric nuclear matter Wang and Li (2013); Brown (2017); Yang and Piekarewicz (2018); Brown et al. (2020); Xu et al. (2022); Pineda et al. (2021); Novario et al. (2023); Huang et al. (2023); An et al. (2023); Bano et al. (2023); König et al. (2024). With the improved measurements, much more data of charge radii of exotic nuclei far away from the $\beta$-stability line have been compiled Angeli and Marinova (2013); Li et al. (2021); Day Goodacre et al. (2021); Yang et al. (2023). The distinguished aspects of nuclear charge radii, involving the shell closures and odd-even staggering (OES) effects, are generally observed throughout the entire nuclide chart Campbell et al. (2016); Garcia Ruiz and Vernon (2020).
@@
The discontinuity changes in charge radii are naturally observed across the traditional neutron magic numbers $N=28$, $50$, $82$, and $126$ Procter et al. (2012); Gorges et al. (2019); Sharma et al. (1993); Goddard et al. (2013); Hammen et al. (2018); de Groote et al. (2020). Many methods pertaining to the effect of shell closure on nuclear charge radii are discussed Wang and Li (2013); Nakada (2019); Perera et al. (2021). As demonstrated in Ref. Sharma et al. (1995), this shell closure effect results from the rather small isospin dependence of the spin-orbit term. Especially striking among these fully filled shells has been found is the intermediate mass region nuclei featuring the $N=28$ neutron-closure shell. The inverted parabolic-like shape between $N=20$ and $N=28$ and odd-even staggering (OES) in charge radii are observed remarkably along Ca isotopes. The same scenario can also be found in K isotopes, but with weakened amplitudes Angeli and Marinova (2013). Across the $N=28$ shell closure, the experimentally observed strong increases in charge radii are naturally manifested for K Koszorús et al. (2019, 2021) and Ca Garcia Ruiz et al. (2016) isotopes, and the odd-even oscillation behaviors have been weakened as well. Recently, charge radii of ${}^{55,56,58-68,70}\mathrm{Ni}$ isotopes have been measured by collinear laser spectroscopy method Malbrunot-Ettenauer et al. (2022); Sommer et al. (2022). It suggests that the trend of charge radii across the $N=28$ shell closure is similar to Ca isotopes. Between neutron numbers $N=28$ and $N=40$, the nuclear charge radii exhibit a universal pattern that is independent of atomic number Kortelainen et al. (2022). These local variations pose great challenges to our knowledge of understanding nuclear force.
@@
The undertaken efforts have been devoted to describing the global trend of charge radii in nuclear chart. The sophisticated algebraic expression involving the Casten factor $P$ can also reproduce the shell effects Casten et al. (1987); Angeli (1991). Subsequently, a five-parameter formula has been proposed by introducing the OES effects Sheng et al. (2015), in which neutron-proton ($np$) interaction originating from the valence proton and neutron contributes to the local variation of nuclear size. This indicates that the $np$ correlation plays an essential role in accessing the fine structure of nuclear size. Miller, $et~{}al.$ point out that short-range neutron-proton tensor interactions cause the protons to move closer to the outside neutrons of a nucleus, thereby increasing the charge radius Miller et al. (2019a). This emphasizes that $np$ correlation results in the changes of the calculated root-mean-square (rms) charge radius. This short-range correlations (SRCs) imply a decrease in the occupation of the states below the Fermi level and a partial occupation in the states above it Souza et al. (2020). Recent study further verifies that the nuclear SRCs induce the high-momentum fluctuations in the nuclear medium, and the role of $np$-correlation has a non-negligible influence on nuclear size Cosyn and Ryckebusch (2021).
@@
As demonstrated in Ref. Zawischa (1985), there is an attractive interaction between proton and neutron pairs, the origin of OES behaviors in nuclear charge radii is due to the four-particle correlations (or being $\alpha$ cluster). The relativistic density functionals cannot describe the fine structure of charge radii along calcium isotopic chain well Geng et al. (2003); Blaum et al. (2008); Xia et al. (2018); Zhang et al. (2022). To reproduce the OES and shell effect of charge radii along calcium isotopes, a modified relativistic mean field plus Bardeen-Cooper-Schrieff (BCS) equation ansatz, namely RMF(BCS)*, has been proposed An et al. (2020). In which the semi-microscopic correction originating from the neutron and proton pairs condensation is introduced by solving the state-dependent BCS equations. The calculated results are in good agreement with the data and compatible with the sophisticated Fayans energy density functional (EDF) approach where the surface pairing dominates a role in reproducing the local variation of charge radii Reinhard and Nazarewicz (2017); Miller et al. (2019b). However, for odd-$Z$ isotopic chain, the OES effects in charge radii are overestimated by RMF(BCS)* approach An et al. (2022a). This may result from the fact that the correlation between the unpaired neutron and proton due to the violation of time-reversal symmetry is ignored. The same scenarios can also be encountered in Fayans EDF model where the OES effects in charge radii are overestimated for K Koszorús et al. (2021) and Cu de Groote et al. (2020) isotopes as well. Obviously, the influence coming from the unpaired nucleons is still excluded out the pairing spaces due to the blocking approximation.
@@
So far, the existing theoretical models can hardly give a unified description of charge radii around $N=28$ mass region. Hence the required model should be able to describe simultaneously the shell effect and OES in nuclear charge radii. In this work, the charge radii are calculated along $Z=19$-$29$ isotopic chains that include the nuclei featuring the $N=28$ magic shell and the new magicity of $N=32$ and $34$. In order to reduce the overestimated OES behaviors of charge radii along odd-$Z$ isotopic chains, the $np$ correlation originating from the simultaneously unpaired neutron and proton is taken into account in the improved charge radii formula. Generally, the gradually shrunken trend of nuclear charge radii is observed around the traditional neutron magic numbers along a long isotopic chain Angeli and Marinova (2013); Li et al. (2021). One should also mention the great interest for the properties of the neutron number $N=32$ and $34$ isotopes that have been proposed as new-magicity nuclei in certain isotopic chains. As demonstrated in Ref. Reiter et al. (2018), the new magicity of $N=32$ is vanished beyond $Z=22$. Meanwhile, the sub-shell closure effect of $N=34$ is weakened across $Z=20$ proton number Crawford et al. (2010); Steppenbeck et al. (2017); Meisel et al. (2020). The sensitive indicators of new-magicity associating with the evolution of nuclear charge radii are also paid more attention along $N=32$ and $34$ isotonic chains in our calculations.
@@ II Theoretical framework
The relativistic mean-field (RMF) theory with different parametric versions has made considerable success in describing various phenomena in nuclear physics Meng and Ring (1996); Vretenar et al. (2005); Liang et al. (2015); Meng (2016); Cao and Ma (2004); Paar et al. (2003, 2004); Zhou et al. (2003); Meng et al. (2006); Zhou et al. (2010); Li et al. (2012); Xu et al. (2015); Niks̆ić et al. (2014); Sun et al. (2018); Zhang et al. (2020, 2021); Rong (2023). In this work, the standard non-linear self-coupling Lagrangian density is employed Ring et al. (1997). The Dirac equations with effective fields $S(\mathbf{r})$ and $V(\mathbf{r})$ are derived through variational principle. To obtain the ground-state properties of finite nuclei, the quadrupole deformation parameter $\beta_{20}$ becomes constrained in the self-consistent iterative process. Therefore, the Hamiltonian formalism can be recalled as follows:
@@
where $M$ is the mass of nucleon, $\lambda$ is the spring constant, and $\mathbf{Q}$ is the intrinsic quadrupole moment. Those values of $\beta_{20}$ are changed from $-0.50$ to $0.50$ with the interval range of $0.01$.
Recent study shows that it is essential to include the staggering behaviors of charge radii in validation protocol Dong et al. (2022, 2023). As mentioned above, a new ansatz has been proposed for describing the OES and shell closure effects in the charge radii of calcium isotopes An et al. (2020). This modified formula can also reproduce the global trend of charge radii along various even-$Z$ isotopic chains An et al. (2022b). Although the shell effects and OES in nuclear charge radii are remarkably described well, the validation protocol is just considered by the chosen nuclei 44Ca and 126Sn An et al. (2020). Besides, the $np$ correlation originating from the simultaneously unpaired proton and neutron is excluded; this leads to the overestimated OES of charge radii for odd-$Z$ isotopic chains An et al. (2022a, c). Therefore, the improved version should be proposed, especially the $np$ correlation coming from the unpaired nucleons must be considered. To account for the implications of the observed odd-even oscillations of charge radii along odd-$Z$ isotopic chains, the further modified mean-square charge radii formula is proposed as follows (in units of fm2),
@@
The first term represents the charge distribution of point-like protons and the second term is due to the finite size of protons Gambhir et al. (1990). Here, the quantity of the proton radius takes the values about $0.84$ fm Tiesinga et al. (2021); Lin et al. (2022). As shown in Eq. (2), the quantity of $\mathcal{D}$ is associated with the Cooper pairs condensation Dussel et al. (2007). Furthermore, this quantity can be used to measure the Fermi surface diffuseness encoded by various eigenfunctions Mang et al. (1965). Its value is calculated self-consistently by solving the state-dependent BCS equations Geng et al. (2003); An et al. (2018). The difference of $\mathcal{D}$ coming from neutrons and protons has been attributed to the $np$ correlation around Fermi surface An et al. (2020, 2022b).
It is mentioned that the amplitude of the OES of charge radii in potassium isotopes is obviously lower than the case of calcium isotopes due to the last unpaired proton Angeli and Marinova (2013); Li et al. (2021). For the theoretical studies, the developed RMF(BCS)* approach and Fayans EDF model can describe the OES of charge radii along even-$Z$ isotopic chains well An et al. (2020); Reinhard and Nazarewicz (2017). However, along odd-$Z$ isotopic chains, the overestimated OES in charge radii is obtained by both of the RMF(BCS)* approach and Fayans EDF model de Groote et al. (2020); An et al. (2022a); Koszorús et al. (2021). The correlation between the simultaneously unpaired proton and neutron may melt this tension. The last term in this expression just represents the $np$ correlation deriving from the simultaneously unpaired neutron and proton. This means the quantity of $\delta$ equals to zero for even-even, odd-even, and even-odd nuclei. The values of $a_{0}=0.561$ and $\delta=0.355$ are calibrated by fitting the odd-even oscillation and the inverted parabolic-like shape of charge radii along potassium and calcium isotopes under effective force NL3. And then the extrapolated calculations are performed for $Z=21$-$29$ isotopic chains. The PK1 set gives a more reasonable baryonic saturation density and a reasonable description of spin-orbit splittings and single-particle energies Long et al. (2004). In addition, NL3 parameter set improves the description of the ground-state properties of many nuclei and simultaneously provides an excellent description of the collective excited states properties in finite nuclei Lalazissis et al. (2009). Thus these two effective forces are also employed to access the global evolution of charge radii along $Z=19$-$29$ isotopic chains. It should be particularly interesting for quantitatively understanding the reduced OES of charge radii along odd-$Z$ isotopic chains and the trend of changes of nuclear charge radii across the $N=28$ shell closure.
@@ III Results and Discussions
In this work, we focus on the systematic behaviors of charge radii along $Z=19$-$29$ isotopic chains. The bulk properties of finite nuclei are calculated by relativistic mean field theory with NL3 Lalazissis et al. (1997), PK1 Long et al. (2004), and NL3 Lalazissis et al. (2009) effective forces. The pairing correlations are treated by the state-dependent Bardeen-Cooper-Schrieffer method, which can capture the ground-state properties of finite nuclei well Geng et al. (2003). The pairing strength is determined through the empirical odd-even mass staggering of calcium isotopes Bender et al. (2000). To reflect the universality of our results, the pairing strength is chosen to $V_{0}=350$ MeV fm3 for NL3 set, but $V_{0}=380$ MeV fm3 for both PK1 and NL3 interaction forces, respectively. Based on the fact that the nuclei with featuring new magicity have been observed at the neutron numbers $N=32$ and $34$. It is instructive to investigate the universal behaviors of charge radii along $N=28$, $32$ and $34$ isotones.
@@ III.1 Shell quenching of nuclear charge radii at $N=28$
The inverted parabolic-like behavior describing the charge radii of the calcium isotopes between $N=20$ and $28$ as well as the amplitude of OES are much more pronounced for the neighboring odd-$Z$ element potassium Touchard et al. (1982). The inverted parabolic-like curves are symmetric to the $1f_{7/2}$ mid-shell neutron number $N=24$ for potassium and calcium isotopic chains Angeli and Marinova (2013). In Fig. 1, the rms charge radii of potassium and calcium isotopes as a function of neutron number are plotted by the relativistic mean-field (RMF) model with effective forces NL3, PK1, and NL3. One can find that the inverted parabolic-like shapes of potassium and calcium isotopes are compared against well by NL3 and NL3 parametrization sets, while the quantitative level of charge radii is systematically underestimated by PK1 force. Beyond $N=28$ shell closure, the sharp increases are identified by these three forces. In Ref. Borzov and Tolokonnikov (2022), the universal increase of the charge radii after crossing the neutron shell at $N=28$ is attributed to the quasiparticle-phonon coupling.
@@
Figure 1: Evolution in rms charge radii versus neutron number for (a) potassium and (b) calcium isotopes. The effective forces NL3 (open circle), PK1 (open square), and NL3 (open triangle) are employed in relativistic mean-field framework. The corresponding odd-even staggerings are also depicted in (c) and (d). Experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021); Koszorús et al. (2021), which are represented by solid diamond with systematic error bands (light blue).
@@
The characteristic aspects in charge radii are the odd-even oscillation behaviors along calcium isotopic. Reference Talmi (1984) demonstrates that the main contribution to the odd-even effect is ascribed to quadrupole and higher even-order core polarization, in which the strong and attractive interaction between the valence neutrons and core protons has been considered. These peculiarities have been ascribed to the changes of dynamic deformation for different orders Barranco and Broglia (1985). Qualitatively the OES of radii as a function of $N$ is associated with the reduction of core polarization due to unpaired nucleons Reehal and Sorensen (1971). This mechanism also provides a straightforward explanation for the odd-even effect as a function of $Z$. Besides, high-order radial moment offers key information on the nuclear charge radius Reinhard et al. (2020).
@@
As shown in Ref. Koszorús et al. (2021), the amplitudes of OES in charge radii for potassium isotopes are overestimated by Fayans EDF. Fayans EDF model points out that surface pairing interactions cause the origin of OES in nuclear charge radii Reinhard and Nazarewicz (2017). However, the time-reversal states are excluded in tackling pairing correlations, especially for simultaneously unpaired neutron and proton in potassium isotopes. The same scenario is encountered in Ref. An et al. (2022a), where the states that violate time-reversal symmetry are tackled by the blocking approximation. This leads to the most overestimated amplitudes of OES of charge radii in potassium isotopes as well. In Ref. An et al. (2020), $np$ correlation originating from the difference of Cooper pair condensation is introduced by solving the state-dependent BCS equations. The modified root-mean-square charge radii formula can describe well the OES and shell effects in charge radii along even-$Z$ isotopic chains. But for odd-$Z$ cases, the OES is overestimated due to the absence of $np$ correlation coming from the simultaneously unpaired neutron and proton An et al. (2022a). As shown in Fig 1, the modified formula can reproduce the OES of charge radii between $N=20$ and $28$ along potassium and calcium isotopes.
@@
In Fig. 1 (a), charge radii of 36-38K isotopes are slightly overestimated by the NL3 and NL3 forces. By contrast, the calculated results for 50K are slightly underestimated by these two forces. The experimental mass measurement suggests that 51K has a relative large shell gap with respect to the neighboring counterparts Rosenbusch et al. (2015). Of high interest is also the experimental determination of the charge radii across the neutron number $N=32$ with emerging the new magicity. A signature of the magic character can be reflected from the sudden increase in charge radii. Charge radii of exotic isotope 52K has been performed by the collinear resonance ionization spectroscopy method Koszorús et al. (2021). This latest measurement suggests that no rapid increase in charge radii can be found across 51K. Beyond 50K, these three effective forces give almost similar trends and show a smooth increase from 51K to 52K.
@@
In Fig. 1 (b), charge radii of 37-39Ca isotopes obtained by PK1 force are compared well against NL3 and NL3 forces. NL3 parameter set gives almost comparable results with respect to NL3 force. With the increasing neutron number occupation at the $p_{3/2}$ orbital, the unexpected linear increase of the charge radius occurs Enciu et al. (2022). This can also be used to understand the emergency of the abrupt increase of charge radii across $N=28$ shell closure as shown in Fig. 1. The neutron number $N=32$ is identified as a new magic number in calcium region An et al. (2018); Wienholtz et al. (2013); Gallant et al. (2012). Therefore, the implication of this experimental value for the charge radius of 53Ca will be urgently expected. As encountered in potassium isotopes, the charge radii beyond $N=34$ show same steep increasing trend with these effective forces.